840 meters in 15 seconds____ meters in 3 seconds
Answer by ptaylor(2198)

840 meters in 15 seconds____ meters in 3 seconds

840 meters is to 15 seconds as x meters is to 3 seconds and this ratio is written as follows:
840/15=x/3 multiply each side by 15 (or cross multiply)
3*840=15x or
2520=15x divide both sides by 15
x=168 meters
CK
3 divides into 15 5 times so:
168*5=840
840=840

Hope this helps---ptaylor